OpenStudy (anonymous):

find the height of a prism of a base if the volume is 9 and the area of the base is 18

OpenStudy (anonymous):

volume of a prism = area of base x height so plug in your 2 values and solve the equation

OpenStudy (anonymous):

i assume that the prism is standing on its end!

OpenStudy (anonymous):

if the base of a right prism is a rhombus with diagonals 10 in and 12 in and the volume of the prism is 60 cu in find the height of the prism please help!

OpenStudy (anonymous):

i really dont understand geometry at alll!

OpenStudy (anonymous):

well we usually talk about the length and cross sectional area of a prism. I prism is a 3 dimensional object with a uniform cross section - a bit like a toblerone bar of chocolate

OpenStudy (anonymous):

ok . could you please explain to me how to solve the problem

OpenStudy (anonymous):

well assume the prism is resting on its base which is a rhombus then its volume = area of the rhombus x the height so 60 = area of rhombus x h so we need to calculate the area of the rhombus

OpenStudy (anonymous):

hold on a bit - to be honest i'm stuck with that bit

OpenStudy (anonymous):

ok so its going to be d1d2 divided by 2 to solve for area

OpenStudy (anonymous):

heck! is that the area of a rhombus - you've got that from a maths book or by other means??

OpenStudy (anonymous):

other means but when i calulate it it comes out to 60 = 60

OpenStudy (anonymous):

i dont remember to be honest

OpenStudy (anonymous):

well d1d2 /2 = 30/2 = 15 so qwe have 60 = 15 x h h = 4 height is 4

OpenStudy (anonymous):

where did u get 15 from?

OpenStudy (anonymous):

where did you get 30 from im lost

OpenStudy (anonymous):

sorry - i must be tired its 12 x 10 divided by 2 giving 60 - either the d1d2 formula is wrong or volume = 60 is wrong. hold on and i'll check my maths stuff

OpenStudy (anonymous):

ok - the volume d1d2/2 is correct so the answer must be h = 1 60 = 60 x h h=1

OpenStudy (anonymous):

is that possible?

OpenStudy (anonymous):

lol - thats wrong because the diagonals are in inches and the volume is in cubic feet right 60 cu feet = 60 x 12^3 cu ins = 103680 cu ins so 103680 = 60 x h h = 12^3 = 1728 ins or 144 ft

OpenStudy (anonymous):

phew - i think thats enough for today my brain is tired

OpenStudy (anonymous):

how did you get to that?

OpenStudy (anonymous):

well the volume is 60 cu ft isn't it

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i 'm going crazy - i've just checked your earlier post - its 60 cu ins there

OpenStudy (anonymous):

in which case h = 1 inch as i said earlier

OpenStudy (anonymous):

that is possible brit

OpenStudy (anonymous):

lol okay thank you but im not done working your brain 2 more questions if you dont mind

OpenStudy (anonymous):

give me a medal first! lol

OpenStudy (anonymous):

ok so the diagonal of a cube is 7\[\sqrt{3}\] find the voulme of the cube

OpenStudy (anonymous):

i assume you mean the diagonal of one side of the cube. by pythagoras (7sqrt3)^2 = 2x^2 (where x = one side of cube) 147 = 2x^2 x = sqrt(147/2) = 8.5732 vlume of cube = (8.5732)^3 = 630.13 correct to 2 dec places

OpenStudy (anonymous):

are u still there